A frame has 50 Bytes. The BER (bit error ratio) of the comunication link is 10^-3. Calculate the probability of these events: (1). The frame errives intact at the destination. and (2). The frame has exactly 2 bit in error.
Answer #1

Solution:

Frame size is= 50 bytes

The BER (bit error ratio) of the communication link is 10^-3.

which means it is = 0.1%

1)

Total number of bits are = 50*8 = 400

the probability that they have arrived intact = 400/1000 = 0.4

2)

So the probability of having exactly two bit in error will be

when two bits are already infected

450-2/1000 = 3988/1000 = 0.398

I hope this helps if you find any problem. Please comment below. Don't forget to give a thumbs up if you liked it. :)